# Spiking Neural Network as Adaptive Event Stream Slicer

### v2

@@ Abstract
Event-based cameras are attracting significant interest as they provide rich edge information, high dynamic range, and high temporal resolution. Many state-of-the-art event-based algorithms rely on splitting the events into fixed groups, resulting in the omission of crucial temporal information, particularly when dealing with diverse motion scenarios (e.g., high/low speed). In this work, we propose SpikeSlicer, a novel-designed plug-and-play event processing method capable of splitting events stream adaptively. SpikeSlicer utilizes a low-energy spiking neural network (SNN) to trigger event slicing. To guide the SNN to fire spikes at optimal time steps, we propose the Spiking Position-aware Loss (SPA-Loss) to modulate the neuron’s state. Additionally, we develop a Feedback-Update training strategy that refines the slicing decisions using feedback from the downstream artificial neural network (ANN). Extensive experiments demonstrate that our method yields significant performance improvements in event-based object tracking and recognition. Notably, SpikeSlicer provides a brand-new SNN-ANN cooperation paradigm, where the SNN acts as an efficient, low-energy data processor to assist the ANN in improving downstream performance, injecting new perspectives and potential avenues of exploration.
@@ 1 Introduction
  • Extensive experiments demonstrate that SpikeSlicer significantly improves model performance by up to 22.3% in object tracking and 19.2% in recognition with only a 0.32% increase in energy consumption.
@@ 4.2 Expert’s Arena: Mastering Adaptive Event Slicing with SNN-ANN Collaboration
Event-based Object Tracking. Since the tracking task is highly sensitive to temporal information, dynamic event slicing is of great importance. Tab. 1 shows that the tracking performances under the SpikeSlicer have a significant improvement in terms of representative success rate (RSR), representative precision rate (RPR), and overlap precision (OP). For instance, TransT ’s performance on OP0.50 and OP0.75 improved by 9.8% and 27.7% compared to its original result under the HDR scenario. When compared to results on the fixed event (the number of the fixed-sliced event is aligned with the number of dynamic-sliced events to ensure a fair comparison), our method achieves favorable gains in the overall RSR, i.e., 62.4 vs. 51.0.
@@
Table 1: Quantitative comparison on FE108. There are four challenging scenarios, including high dynamic range (HDR), low light (LL), fast motion with and without motion blur (FWB & FNB) and all testing datasets (ALL). [method]+SpikeSlicer represents the results based on our adaptive event slicing method and the results of [method] are tested on the original fixed-sliced event dataset from [3]. To ensure a fair comparison, fix event indicates that the model is tested on a dataset of fixed-sliced event frames, where the number of fixed event frames is the same as the number of dynamically sliced event frames by using SpikeSlicer. Best performances are denoted by deep green.
@@
DiMP+SpikeSlicer 52.3 65.5 21.4 78.7 69.4 92.4 43.2 97.5 52.4 64.8 26.6 78.8 62.2 81.0 21.8 97.2 56.0 71.4 23.3 85.1
PrDiMP [32] 44.3 52.8 19.6 66.3 44.6 48.2 8.9 69.5 67.0 89.9 33.6 99.7 60.6 75.8 29.7 93.3 53.0 65.0 23.3 80.5
@@
PrDiMP+SpikeSlicer 53.5 67.2 22.1 79.7 70.1 94.5 41.8 97.4 70.7 88.3 55.5 94.3 56.3 70.5 25.9 83.9 59.2 75.1 29.4 86.7
TransT [27] 54.0 68.5 22.0 81.7 67.9 88.4 40.4 96.7 67.0 92.4 30.5 99.6 58.0 73.6 27.5 87.2 58.1 74.7 25.7 87.7
@@
TransT+SpikeSlicer 57.7 75.2 28.1 82.6 70.7 93.6 42.7 99.0 74.9 97.7 61.1 98.6 58.7 75.6 29.6 84.6 62.4 81.2 34.1 88.9

### v3

@@ Abstract
Event-based cameras are attracting significant interest as they provide rich edge information, high dynamic range, and high temporal resolution. Many state-of-the-art event-based algorithms rely on splitting the events into fixed groups, resulting in the omission of crucial temporal information, particularly when dealing with diverse motion scenarios (e.g., high/low speed). In this work, we propose SpikeSlicer, a novel-designed plug-and-play event processing method capable of splitting events stream adaptively. SpikeSlicer utilizes a low-energy spiking neural network (SNN) to trigger event slicing. To guide the SNN to fire spikes at optimal time steps, we propose the Spiking Position-aware Loss (SPA-Loss) to modulate the neuron’s state. Additionally, we develop a Feedback-Update training strategy that refines the slicing decisions using feedback from the downstream artificial neural network (ANN). Extensive experiments demonstrate that our method yields significant performance improvements in event-based object tracking and recognition. Notably, SpikeSlicer provides a brand-new SNN-ANN cooperation paradigm, where the SNN acts as an efficient, low-energy data processor to assist the ANN in improving downstream performance, injecting new perspectives and potential avenues of exploration. Our code is available at https://github.com/AndyCao1125/SpikeSlicer.
@@ 1 Introduction
  • Extensive experiments demonstrate that SpikeSlicer significantly improves model performance by up to 11.9% in object tracking and 19.2% in recognition with only a 0.32% increase in energy consumption.
@@ 4.2 Expert’s Arena: Mastering Adaptive Event Slicing with SNN-ANN Collaboration
Event-based Object Tracking. Since the tracking task is highly sensitive to temporal information, dynamic event slicing is of great importance. We provide two versions for adaptive slicing: SpikeSlicer-Base (B) and SpikeSlicer-Small (S). The detailed consumption of these two versions are introduced in the ablation study. Tab. 1 shows that the tracking performances under the SpikeSlicer have a significant improvement in terms of representative success rate (RSR), representative precision rate (RPR), and overlap precision (OP). For instance, TransT with SpikeSlicer-S’s performance on OP0.50 and OP0.75 improved by 5.8% and 3.2% compared to its original result under the HDR scenario. When compared to results on the fixed event (the number of the fixed-sliced event is aligned with the number of dynamic-sliced events to ensure a fair comparison), our method achieves favorable gains in the overall RSR, i.e., 63.6 vs. 51.0.
@@
Table 1: Quantitative comparison on FE108. There are four challenging scenarios, including high dynamic range (HDR), low light (LL), fast motion with and without motion blur (FWB & FNB) and all testing datasets (ALL). [method]+SpikeSlicer-B/S represents the results based on our adaptive event slicing method with base (B) or small (S) SNN version. The results of [method] are reproduced on the original fixed-sliced event dataset. To ensure a fair comparison, fix event indicates that the model is tested on a dataset of fixed-sliced event frames, where the number of fixed event frames is the same as the number of dynamically sliced event frames by using SpikeSlicer. Best performances are denoted by deep green.
@@
DiMP+SpikeSlicer-B 53.3 67.3 21.9 79.8 69.8 92.3 47.2 96.5 64.1 83.4 27.9 97.2 54.4 67.4 27.7 81.2 57.3 73.0 25.8 86.1
DiMP+SpikeSlicer-S 56.0 72.0 27.0 80.7 70.0 92.2 50.6 95.0 66.7 86.1 38.4 96.9 56.4 70.6 31.1 81.4 59.6 76.8 30.9 86.4
PrDiMP [32] 50.3 62.2 19.4 77.8 68.8 90.4 41.9 97.0 56.6 68.8 11.2 98.1 53.4 64.7 23.4 82.7 54.5 67.4 20.4 85.8
@@
PrDiMP+SpikeSlicer-B 53.7 67.1 22.2 80.2 70.1 94.5 41.8 97.2 70.7 89.2 54.2 93.9 56.1 70.3 25.7 83.7 59.2 75.3 29.1 86.8
PrDiMP+SpikeSlicer-S 55.2 70.7 24.9 79.9 71.3 95.7 45.1 97.7 72.5 91.4 59.2 95.4 57.6 73.0 27.6 83.8 60.9 78.2 32.3 87.2
TaMOs [33] 37.9 43.5 1.5 66.9 46.4 30.5 0.2 87.8 50.8 54.2 1.0 96.7 40.9 36.2 0.7 77.1 42.5 41.2 1.0 78.8
TaMOs (fixed event) 44.0 57.0 3.9 72.7 49.5 48.0 0.3 90.0 46.2 34.0 0.5 69.7 43.4 46.0 1.5 78.4 45.1 48.5 2.1 77.1
TaMOs+SpikeSlicer-B 40.4 45.4 1.7 72.2 47.1 34.6 0.1 91.7 48.7 45.7 0.3 98.5 43.6 42.0 1.0 81.7 44.2 43.1 1.1 83.3
TaMOs+SpikeSlicer-S 41.4 48.1 1.4 71.7 48.5 36.1 0.3 95.5 45.0 18.7 0.1 98.7 40.7 29.6 0.5 80.4 43.0 36.3 0.8 82.9
TransT [27] 55.8 69.4 27.7 82.0 70.5 90.9 49.7 98.3 74.1 98.5 55.5 99.9 58.6 73.9 30.2 87.2 61.3 78.3 33.8 89.2
@@
TransT+SpikeSlicer-B 58.0 73.4 29.4 83.8 71.9 95.9 47.7 99.3 73.4 95.7 56.9 96.2 60.5 76.6 32.0 88.4 62.1 79.9 34.6 88.7
TransT+SpikeSlicer-S 59.1 75.2 30.9 84.4 72.9 97.9 52.2 99.1 76.6 99.5 65.7 99.8 60.2 76.5 32.3 87.5 63.6 82.3 37.5 90.1
@@ 4.3 Analysis of the Adaptive Slicing Method
Table 7: Ablation studies for different network sizes of the SpikeSlicer. The comparison includes the model consumptions and tracking performances of PrDiMP.
Methods Model Consumptions Performances
Params (M) OPs (G) Energy (mJ) RSR OP.50 OP.75 RPR
SpikeSlicer-Base 45.11 0.73 0.85 59.24 75.25 29.12 86.82
SpikeSlicer-Small 0.42 0.56 0.69 60.88 78.19 32.34 87.19
@@ 4.4 Ablation Study
Ablation Study on Different Network Sizes of the SpikeSlicer. To evaluate the performance of SpikeSlicer under different network sizes, we conduct an additional ablation study (in Tab. 7) with a smaller variant, SpikeSlicer-Small. This lightweight model contains only 0.42M parameters—significantly fewer than the base model’s while achieving comparable or better performance across key metrics. This compact design demonstrates its potential for efficient deployment on hardware platforms, providing a strong foundation for real-world applications requiring lightweight neural networks.